6. (a) Draw the logic circuit for the following expression: AB + A(B+C)

(b) Simplify the expression in 6(a) by using the rules of Boolean algebra provided in (5). (c) Draw the simplified logic gate circuit derived in (b)


1
Expert's answer
2021-11-24T17:01:46-0500

(a)



(b) "AB+A(B+C)=AB+AB+AC=AB+AC=A(B+C)"

(c)